OpenStudy (afifarahman):

prove that right bisector of the chord joining the points A(2,2) and B(3,1) of the circle x^2+y^2-4x-2y+4=0 pass through centre of the circle.

OpenStudy (anonymous):

za maray :P

OpenStudy (science0229):

x^2+y^2-4x-2y+4=0 is the same as (x-2)^2+(y-1)^2=1 The midpoint of AB is (5/2, 3/2) The center point of the circle is (2,1) the equation of a line passing through those 2 points is y=x-1

OpenStudy (science0229):

The equation of the line passing through A and B is y=-x+4 Since those 2 equations are perpendicular, the proof if complete.

OpenStudy (science0229):